OpenStudy (anonymous):

How many possible number of solutions does this equation have? x + 3y = 2 2x + 6y = -3

OpenStudy (phi):

can you multiply the first equation by 2 (that means both sides, all terms) ?

OpenStudy (anonymous):

no...

OpenStudy (phi):

multiply each term: the x, the 3y and the 2 (on the other side of the =) by 2 to make a new equation

OpenStudy (phi):

you should get 2x+6y= 4

OpenStudy (phi):

now you have two equations 2x+6y= 4 2x + 6y = -3 notice that 2x+6y has to be 4 and *at the same time* -3

OpenStudy (phi):

or if you subtract the two equations, you get 0 = 4-(-3) or 0 = 7 which is not true. That is a hint there is no solution

OpenStudy (anonymous):

perfect, i understand it now. thanks (: